OpenStudy (anonymous):

what is the value of 3x-4 if 2x+1=5

OpenStudy (anonymous):

2x+1=5 x=2 3(2)-4=0

OpenStudy (anonymous):

so its 2

OpenStudy (anonymous):

x=2 but the problem 3x-4=0

OpenStudy (anonymous):

thats not right its multiple chicie and thats not a choice a. 2 b. 4 c.6 d. 8

OpenStudy (anonymous):

oh lol my bad it is 2 lol sorry i was useing 4 insted of 6 :) i apologize